A hydrogen atom moving with a speed 'v' absorbs a photon of wavelength 127nm and stops.Then the value of 'v' is nearly...........


 


Pls explain.......

Grade:12

2 Answers

aman gupta
13 years ago

energy of a photon is = hc/λ

visible light lies in 390-750 nm 

maximum energy for 390 nm

h=6.626 x  10-34 

c= 3 x 108

λ=390 x 10-9

 

E=5.09 x 10-19 Joules

Geoffrey Richards
13 years ago

The ans comes out to be v=3.1 m/s

 

Use the formula     lambda = h/mv

lambda = 127nm

h = 6.626*10-34 Js

mass (m)= 1.67*10-27 kg

 

Now evaluate do it as a practice ...

U get v =3.1 m/s